OpenStudy (anonymous):

Evaluate -6 - (-1). -7 5 -5 6

OpenStudy (anonymous):

@misty1212 , @welshfella

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

When both numbers are negative and you are subtracting you basically add. So -6 - (-1) = what

OpenStudy (welshfella):

-6 - (-1) double negative = a + :- = -6 + 1 = ?

OpenStudy (anonymous):

-6 + 1 = 5 , and I think I have to keep the sign of the larger number right?

OpenStudy (anonymous):

No you keep it negative

OpenStudy (anonymous):

No, I mean for the final answer, doesn't the final answer have to be negative since -6 is the larger absolute value?

OpenStudy (anonymous):

Yes, the final answer is negative

OpenStudy (welshfella):

its illustrated on a number line